OpenStudy (anonymous):

Which expression is equivalent to -3 – (7.5 + 4)? A. - (7.5 + 4) + 3 B. (7.5 + 4) - 3 C. - (3 - 7.5) + 4 D. - (7.5 + 4) - 3 I will fan im super in need of some help

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

it is the commutative law in action

OpenStudy (anonymous):

hey!!!

OpenStudy (misty1212):

\[a+b=b+a\] so \[-3-(7.5+4)=-(7.5+4)-3\]

OpenStudy (ahvei):

B

OpenStudy (misty1212):

definitely not B

OpenStudy (anonymous):

I have a couple more qustions